# Normal forms for ordinary differential operators, I

### v1

@@ 3 Normal forms for commuting operators
A self contained proof of this theorem see e.g. in [43, 10.3].
Let $C$ be a projective curve over $K$ and $p\in C$ be a regular $K$-point. Then $C_{0}:=C\backslash p$ is an affine curve. According to theorem 3.1 for any torsion free rank one sheaf ${{\cal F}}$ on $C$ there exists a normalised elliptic ring of differential operators $B\subset D_{1}$ defined uniquely up to a scale transform, such that $B$ is isomorphic to the ring of regular functions on $C_{0}$, $B\simeq{{\cal O}}_{C_{0}}(C_{0})$. Vice versa, any such ring is isomorphic to the ring of regular functions on some affine curve $C_{0}$, which can be compactified with the help of one regular $K$-point.
@@
We’ll call an affine curve $C_{0}$ over $K$ as affine spectral curve if it can be compactified with the help of one regular $K$-point, i.e. if there exists a projective curve $C$ over $K$ and a regular $K$-point $p$ such that $C_{0}\simeq C\backslash p$ (note that such $C$ is uniquely defined up to an isomorphism, see e.g. [12, Ch1., §6]).
@@
defined for any coherent torsion free sheaf ${{\cal F}}$ and any trivialisations $\pi,\hat{\phi}$ (for details see [43, Ch.10]). If ${{\cal F}}={{\cal O}}_{C}$, then $W=A$ and the rank of $A$ is $1$. Moreover, elements of $A$ are Laurent series expansions of functions from ${{\cal O}}_{C}(C\backslash p)$, s.t. the orders of elements of $A$ are the pole orders of corresponding functions. Differential operators corresponding to elements of $A$ via the correspondence from theorems 3.1, 3.2 have the same order as the elements.
@@
Let $B\subset D_{1}$ be an elliptic commutative subring of ODOs. Let $P_{1},\ldots,P_{m}$ be its monic generators over $K$ (any generators have constant highest coefficients, cf. [43, Ch.3]), such that $\mathop{\rm\bf ord}(P_{1})=q$ is coprime with the orders of $P_{2},\ldots,P_{m}$.
@@
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal, $f_{i}\in K[T_{1},\ldots,T_{m}]$. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$. From lemma 3.4 we immediately get that the coefficients of all other normal forms $P_{i}^{\prime}$ are uniquely determined. Obviously, $f_{i}(P_{1}^{\prime},\ldots,P_{m}^{\prime})=0$ for all $i$, and therefore define a set of equations on the coefficients of $\hat{\Phi}(P_{i}^{\prime})\in\mathfrak{B}$ in an affine space. Note that any point of the affine algebraic set determined by these equations defines a set of coefficients of commuting operators from $C(\partial^{q})$, and these operators generate a ring isomorphic to $B$101010with the isomorphism sending them to $T_{1},\ldots,T_{m}$.
@@
We’ll call a commutative ring $B^{\prime}\in C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ generated over $\tilde{K}$ by monic operators $P_{1}^{\prime}=\partial^{q},P_{2}^{\prime},\ldots,P_{m}^{\prime}$, where $P_{2}^{\prime}$ is normalised and $q$ is coprime with the orders of $P_{i}^{\prime}$, $i\geq 2$, as a normalised normal form with respect to a (ordered) set of generators $P_{1}^{\prime},\ldots,P_{n}^{\prime}$, and we’ll call the coefficients of the operators $P_{i}^{\prime}$ as coordinates of $B^{\prime}$ (cf. lemma 3.4).
@@
Then $a:=S^{-1}P^{\prime}S\in\tilde{K}((\tilde{D}^{-1}))$, where $S$ is an operator from lemma 3.3, and $a$ is the uniquely defined monic element in $\tilde{K}((\tilde{D}^{-1}))$, satisfying the equation $f(a,\tilde{D}^{q})=0$.
The proof is obvious in view of remark 3.5 and corollary 3.1.
@@
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal and the order of $w_{1}$ is coprime with the order of $w_{i}$, $i\geq 2$, and the images of $w_{i}$ under the Krichever map (after some choice of $\pi$, $\hat{\phi}$) are monic (cf. remark 3.5).
Then there exist normalised normal forms $B^{\prime}\simeq{{\cal O}}_{C}(C_{0})$ with respect to the ordered set of generators $P_{1}^{\prime}=\partial^{q},\ldots,P_{m}^{\prime}$, where $\mathop{\rm\bf ord}(P_{i}^{\prime})=\mathop{\rm ord}\nolimits(w_{1})$ for all $i\geq 1$, and there is a one to one correspondence between closed points of the affine algebraic set $X_{[{{\cal O}}_{C}(C_{0})]}$ and isomorphism classes of torsion free rank one sheaves ${{\cal F}}$ on $C$ with vanishing cohomologies $H^{0}(C,{{\cal F}})=H^{1}(C,{{\cal F}})=0$.
@@
Let ${{\cal F}}$ be a torsion free rank one sheaf on $C$ with vanishing cohomologies. By theorem 3.1 the triple $(C,p,{{\cal F}})$ corresponds to uniquely defined normalised commutative elliptic subring $B\subset D_{1}$ up to a scale transform. Besides, the generators $w_{i}$ corresponds to formally elliptic differential operators $P_{1},\ldots,P_{m}$ of the same orders, and there exists a scale transform that makes them monic. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$, i.e. $B^{\prime}$ is a normalised normal form w.r.t. $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$ or $P_{i}$. Note that the scale transform is compatible with conjugation by $S$ and that the coefficients of $\hat{\Phi}(P_{i}^{\prime})$ are invariant under any scale transform of $P_{i}^{\prime}$ for all $i$. Thus, ${{\cal F}}$ determines a closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$.
Vice versa, any closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$ determines a normalised normal form $B^{\prime}$ w.r.t. some $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$. By corollary 3.1 there exists an operator $S\in E_{q}$ such that $A^{\prime}:=S^{-1}B^{\prime}S\in\tilde{K}((\tilde{D}^{-1}))$. Note that $W^{\prime}:=F^{\prime}\cdot S$ has support equal to $F^{\prime}$, because $S$ is a monic invertible operator. So, $(W^{\prime},A^{\prime})$ form a Schur pair of rank one. By theorems 3.2 and 3.1 this Schur pair determines a normalised commutative subring $B\subset D_{1}$ of rank one and a torsion free sheaf ${{\cal F}}\simeq\mathop{\rm Proj}\tilde{W^{\prime}}$ of rank one with vanishing cohomologies.
As it was noticed above, the modules $F$, $F^{\prime}$ and $W^{\prime}$ are isomorphic as filtered modules, and all correspondences are compatible with the scale transform. By this reason the maps $p\in X_{[{{\cal O}}_{C}(C_{0})]}\mapsto{{\cal F}}$ and ${{\cal F}}\mapsto p\in X_{[{{\cal O}}_{C}(C_{0})]}$ are mutually inverse. ∎
@@
$P^{\prime}=4\partial^{3}+2u_{0}A_{2,1}\partial+u_{1}A_{1}+\frac{2u_{0}^{2}+u_{2}}{2}(-1+A_{2,1})\int.$
@@
$\tilde{P}=\psi\circ\hat{\Phi}(P)=\begin{pmatrix}&4\\ 4D^{2}&\end{pmatrix}D^{2}+\begin{pmatrix}u_{1}&2u_{0}\\ -2u_{0}D^{2}&-u_{1}\end{pmatrix}+\begin{pmatrix}0&0\\ (-2u_{0}^{2}-u_{2})D^{2}&0\end{pmatrix}D^{-2}$
@@ 4.1 A Newton Region of operators with the property $A_{q}(k)$
Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$, $\mathop{\rm\bf ord}(P)=p$, satisfies condition $A_{k}(0)$ and $Sdeg_{A}(P_{\mathop{\rm\bf ord}(P)-i})=i-1$ for all $i>0$ (such condition holds for an operator $P^{\prime}$ from corollary 2.4, which comes from a generic pair of operators $P,Q\in D_{1}$).
@@ 4.2 One combinatorial lemma
$[D,T_{s-1,j-1,k-1}]=\sum_{t_{1}+\ldots+t_{m}=j-1\atop m=s-j}\binom{k-1}{s-1}g(t_{1},\ldots,t_{m})(L^{(t_{1}+1,\ldots,t_{m})}+\ldots+L^{(t_{1},\ldots,t_{m}+1)})=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(t_{1}^{\prime}-1,t_{2}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(t_{1}^{\prime},\ldots,t_{m-1}^{\prime},t_{m}^{\prime}-1))L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k-1}{s-1}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}$
@@
$T_{s,j,k}=\sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s}g(t_{1}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k-1}{s-1}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}g(t_{2}^{\prime},t_{3}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k}{s}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k}{s}g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k}{s}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}$
@@
$T_{k,j,k}=\sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}g(t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}+\sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=k-j}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}$

### v3

@@ 3 Normal forms for commuting operators
A self contained proof of this theorem see e.g. in [47, 10.3].
Let $C$ be a projective curve over $K$ and $p\in C$ be a regular $K$-point. Then $C_{0}:=C\backslash p$ is an affine curve. According to theorem 3.1 for any torsion free rank one sheaf ${{\cal F}}$ on $C$ there exists a normalised commutative elliptic ring of differential operators $B\subset D_{1}$ defined uniquely up to a scale transform, such that $B$ is isomorphic to the ring of regular functions on $C_{0}$, $B\simeq{{\cal O}}_{C_{0}}(C_{0})$. Vice versa, any such ring $B$ is isomorphic to the ring of regular functions on some affine curve $C_{0}$, which can be compactified with the help of one regular $K$-point.
@@
We’ll call an affine curve $C_{0}$ over $K$ as affine spectral curve if it can be compactified with the help of one regular $K$-point, i.e. if there exists a projective curve $C$ over $K$ and a regular $K$-point $p$ such that $C_{0}\simeq C\backslash p$ (note that such $C$ is uniquely defined up to an isomorphism, see e.g. [16, Ch1., §6]).
@@
defined for any coherent torsion free sheaf ${{\cal F}}$ and any trivialisations $\pi,\hat{\phi}$ (for details see [47, Ch.10]). If ${{\cal F}}={{\cal O}}_{C}$, then $W=A$ and the rank of $A$ is $1$. Moreover, elements of $A$ are Laurent series expansions of functions from ${{\cal O}}_{C}(C\backslash p)$, s.t. the orders of elements of $A$ are the pole orders of corresponding functions. Differential operators corresponding to elements of $A$ via the correspondence from theorems 3.1, 3.2 have the same order as the elements.
If $\mathop{\rm rk}({{\cal F}})=1$, then $\mathop{\rm rk}(A)=1$ and the data $(C,p,{{\cal F}},\pi)$ uniquely determines the class $[B]$ from theorem 3.1, see [47, Rem. 10.24].
@@
Let $B\subset D_{1}$ be an elliptic commutative subring of ODOs. Let $P_{1},\ldots,P_{m}$ be its monic generators over $K$ (any generators have constant highest coefficients, cf. [47, Ch.3]), such that $\mathop{\rm\bf ord}(P_{1})=q$ is coprime with the order of $P_{2}$.
@@
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal, $f_{i}\in K[T_{1},\ldots,T_{m}]$. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$. From lemma 3.4 we immediately get that the coefficients of all other normal forms $P_{i}^{\prime}$ are uniquely determined. Obviously, $f_{i}(P_{1}^{\prime},\ldots,P_{m}^{\prime})=0$ for all $i$, and therefore define a set of equations on the coefficients of $\hat{\Phi}(P_{i}^{\prime})\in\mathfrak{B}$ in an affine space. Note that any point of the affine algebraic set determined by these equations together with equations coming from commutation relations between $P_{2}^{\prime},\ldots,P_{m}^{\prime}$ defines a set of coefficients of commuting operators from $C(\partial^{q})$, and these operators generate a ring isomorphic to $B$888with the isomorphism sending them to $T_{1},\ldots,T_{m}$.
@@
We’ll call a commutative ring $B^{\prime}\in C(\partial^{q})\subset\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$ generated over $\tilde{K}$ by monic operators $P_{1}^{\prime}=\partial^{q},P_{2}^{\prime},\ldots,P_{m}^{\prime}$, where $P_{2}^{\prime}$ is normalised and $q$ is coprime with the order of $P_{2}^{\prime}$, as a normalised normal form of rank one with respect to a (ordered) set of generators $P_{1}^{\prime},\ldots,P_{m}^{\prime}$, and we’ll call the coefficients of the operators $P_{i}^{\prime}$ as coordinates of $B^{\prime}$ (cf. lemma 3.4).
@@
Then $a:=S^{-1}\hat{\Phi}(P^{\prime})S\in\tilde{K}((\tilde{D}^{-1}))$, where $S$ is an operator from lemma 3.3, and $a$ is the uniquely defined monic element in $\tilde{K}((\tilde{D}^{-1}))$, satisfying the equation $f(a,\tilde{D}^{q})=0$.
The proof is obvious in view of remark 3.7 and corollary 3.1. In particular, from theorem below it follows that any monic operator $P^{\prime}$ from lemma together with $\partial^{q}$ form a normal from of some pair of differential operators.
@@
where $I=(f_{1},\ldots,f_{k})$ is a prime ideal and the order of $w_{1}$ is coprime with the order of $w_{2}$, and the images of $w_{i}$ under the Krichever map (after some choice of $\pi$, $\hat{\phi}$) are monic (cf. remark 3.7).
Then there exist normalised normal forms of rank one $B^{\prime}\simeq{{\cal O}}_{C}(C_{0})$ with respect to the ordered set of generators $P_{1}^{\prime}=\partial^{q},\ldots,P_{m}^{\prime}$, where $\mathop{\rm\bf ord}(P_{i}^{\prime})=\mathop{\rm ord}\nolimits(w_{i})$ for all $i\geq 1$, and there is a one to one correspondence between closed points of the affine algebraic set $X_{[{{\cal O}}_{C}(C_{0})]}$ and isomorphism classes of torsion free rank one sheaves ${{\cal F}}$ on $C$ with vanishing cohomologies $H^{0}(C,{{\cal F}})=H^{1}(C,{{\cal F}})=0$.
@@
Let ${{\cal F}}$ be a torsion free rank one sheaf on $C$ with vanishing cohomologies. By theorem 3.1 the triple $(C,p,{{\cal F}})$ corresponds to uniquely defined normalised [^9] Each equivalence class $[B]$ contains a normalised representative, which determines the whole class uniquely. For example, we can choose a monic differential operator of minimal positive order in $B$, and normalise it (conjugating by an appropriate function $f\in\hat{R}$). commutative elliptic subring $B\subset D_{1}$ up to a scale transform. Besides, the generators $w_{i}$ corresponds to formally elliptic differential operators $P_{1},\ldots,P_{m}$ of the same orders, and there exists a scale transform that makes them monic. By proposition 2.3 and lemma 3.4 there exists a uniquely determined Schur operator $S\in\hat{D}_{1}^{sym}$ such that $B^{\prime}:=S^{-1}BS\in C(\partial^{q})$ and $P_{2}^{\prime}=S^{-1}P_{2}S$ is a normalised normal form of $P_{2}$ with respect to $P_{1}$, i.e. $B^{\prime}$ is a normalised normal form w.r.t. $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$ or $P_{i}$. Note that the scale transform is compatible with conjugation by $S$ and that the coefficients of $\hat{\Phi}(P_{i}^{\prime})$ are invariant under any scale transform of $P_{i}^{\prime}$ for all $i$. Thus, ${{\cal F}}$ determines a closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$.
Vice versa, any closed point of $X_{[{{\cal O}}_{C}(C_{0})]}$ determines a normalised normal form $B^{\prime}$ w.r.t. some $P_{1}^{\prime},\ldots,P_{m}^{\prime}$ which have the same orders as $w_{i}$. By corollary 3.1 and remark 3.5 there exists an operator $S\in E_{q}$ such that $A^{\prime}:=S^{-1}\hat{\Phi}(B^{\prime})S\in K((\tilde{D}^{-1}))$. Note that $W^{\prime\prime}:=F^{\prime\prime}\cdot S$ has support equal to $F^{\prime\prime}$, because $S$ is a monic invertible operator of order zero. So, $(W^{\prime\prime},A^{\prime})$ form an embedded Schur pair of rank one [^10] Since $\hat{\Phi}(C(\partial^{p}))$ consists of Laurent polynomials whose coefficients at negative powers of $\tilde{D}$ are not constants (see remark 3.4), and since $S$ is monic and $A^{\prime}\subset K((\tilde{D}^{-1}))$, it follows that $A^{\prime}\cap K[[\tilde{D}^{-1}]]=K$.. By theorems 3.2 and 3.1 this Schur pair determines a normalised commutative subring $B\subset D_{1}$ of rank one and a torsion free sheaf ${{\cal F}}\simeq\mathop{\rm Proj}\tilde{W^{\prime\prime}}$ of rank one with vanishing cohomologies.
As it was noticed above, the modules $F$, $F^{\prime\prime}$ and $W^{\prime\prime}$ are isomorphic as filtered modules, and all correspondences are compatible with the scale transform. By this reason the maps $p\in X_{[{{\cal O}}_{C}(C_{0})]}\mapsto{{\cal F}}$ and ${{\cal F}}\mapsto p\in X_{[{{\cal O}}_{C}(C_{0})]}$ are mutually inverse. ∎
@@
$P^{\prime}=4\partial^{3}+2u_{0}A_{1}\partial+u_{1}A_{1}+\frac{2u_{0}^{2}+u_{2}}{2}(-1+A_{1})\int.$
@@
$\tilde{P}=\psi\circ\hat{\Phi}(P)=\left(\begin{array}[]{cc}u_{1}&2\left(2\tilde{D}^{2}+u_{0}\right)\\ 4\tilde{D}^{4}-2\tilde{D}^{2}u_{0}-2u_{0}^{2}-u_{2}&-u_{1}\\ \end{array}\right)$
@@
Example 3.2.
On the other hand, if we have two operators $L_{2},L_{3}\in D_{1}$ satisfying the equation $L_{3}^{2}=4L_{2}^{3}-g_{2}L_{2}-g_{3}$, where $g_{2},g_{3}$ are two constant coefficients in $K$ (may assume $L_{2}$ is normalized), then apparently $[L_{2},L_{3}]=0$ (cf. [7]).
Consider the normal form $L_{3}^{\prime}$ of $L_{3}$ with respect to $L_{2}$. We fix $k=2,\xi=-1$ in this example, according to Lemma 2.4, $L_{3}^{\prime}$ should be in the form of (within $Hcpc(2)$):
$\hat{\Phi}(L_{3}^{\prime})=2\tilde{D}^{3}+(c_{1,0}+c_{1,1}A_{1})\tilde{D}+(c_{0,0}+c_{0,1}A_{1})+(c_{-1,0}+c_{-1,1}A_{1})\tilde{D}^{-1}$
with $c_{-1,0}+c_{-1,1}=0$.
The calculations of $L_{3}^{\prime}$ will be basically separated into the following parts:
  1. 1.
    Written $L_{2}$ into G-form with unknown coefficients $t_{i}$, notice it’s totally free of $B_{j}$(because $L_{2}\in D_{1}$, and by theorem 2.2), then use lemma 2.14:
    $\hat{\Phi}(L_{2})=\tilde{D}^{2}+t_{0}+t_{1}\Gamma_{1}\tilde{D}^{-1}+(t_{2}\Gamma_{1}-t_{2}\Gamma_{2})\tilde{D}^{-2}+\cdots$
    and correspondingly $L_{3}$ with unknown coefficient $m_{j}$
    $\hat{\Phi}(L_{3})=2\tilde{D}^{3}+m_{0}\tilde{D}+m_{1}+m_{2}\Gamma_{1}+(m_{3}\Gamma_{1}+m_{4}\Gamma_{2})\tilde{D}^{-1}+\cdots$
    Substitute the expression of $(L_{2},L_{3})$ into equation $L_{3}^{2}=4L_{2}^{3}-g_{2}L_{2}-g_{3}$, we get the relation that
    $\begin{cases}m_{0}=3t_{0};\\ m_{1}=\frac{3t_{1}}{2};m_{2}=3t_{1};\\ m_{3}=0;m_{4}=\frac{3}{2}(g_{2}-3t_{0}^{2});\end{cases}$
    with the extra relations with parameter $g_{2},g_{3}$ that
    $\begin{cases}t_{2}=\frac{1}{2}(3t_{0}^{2}-g_{2})\\ -2g_{2}t_{0}+4g_{3}+2t_{0}^{3}+t_{1}^{2}=0\end{cases}$
  2. 2.
    Calculate the Schur operator $S$ in the equation $L_{2}S=S\partial^{2}$.
  3. 3.
    Calculate the normal form $L_{3}^{\prime}$ in the equation $L_{3}S=SL_{3}^{\prime}$ (these calculations were made with the help of Wolfram Mathematica package)
we get
$\hat{\Phi}(L_{3}^{\prime})=2\tilde{D}^{3}+t_{0}A_{1}\tilde{D}-\frac{1}{2}t_{1}A_{1}-\frac{1}{4}(1-A_{1})\left(g_{2}-t_{0}^{2}\right)\tilde{D}^{-1};$
Transferring $L_{3}^{\prime}$ into matrix form, we have
$\psi\circ\hat{\Phi}(L_{3}^{\prime})=\left(\begin{array}[]{cc}-\frac{t_{1}}{2}&2\tilde{D}^{2}+t_{0}\\ \frac{1}{2}\left(4\tilde{D}^{4}-2\tilde{D}^{2}t_{0}-g_{2}+t_{0}^{2}\right)&\frac{t_{1}}{2}\\ \end{array}\right)$
@@ 4.1 A Newton Region of operators with the property $A_{q}(k)$
Suppose $P\in\hat{D}_{1}^{sym}\hat{\otimes}_{K}\tilde{K}$, $\mathop{\rm\bf ord}(P)=p$, satisfies condition $A_{k}(0)$ and
$Sdeg_{A}(P_{\mathop{\rm\bf ord}(P)-i})=i-1$ for all $i>0$ (such condition holds for an operator $P^{\prime}$ from corollary 2.4, which comes from a generic pair of operators $P,Q\in D_{1}$).
@@ 4.2 One combinatorial lemma
$[D,T_{s-1,j-1,k-1}]=\sum_{t_{1}+\ldots+t_{m}=j-1\atop m=s-j}\binom{k-1}{s-1}g(t_{1},\ldots,t_{m})(L^{(t_{1}+1,\ldots,t_{m})}+\ldots+L^{(t_{1},\ldots,t_{m}+1)})=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(t_{1}^{\prime}-1,t_{2}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(t_{1}^{\prime},\ldots,t_{m-1}^{\prime},t_{m}^{\prime}-1))L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k-1}{s-1}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+\\ g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}$
@@
$T_{s,j,k}=\sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s}g(t_{1}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k-1}{s-1}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k-1}{s-1}g(t_{2}^{\prime},t_{3}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=s-j}\binom{k}{s}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k}{s}g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\\ \sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=s-j}\binom{k}{s}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}$
@@
$T_{k,j,k}=\sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}g(t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}+\sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop t_{1}^{\prime}\geq 1,m=k-j}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}+\\ \sum_{t_{2}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}(g(0,t_{2}^{\prime}-1,t_{3}^{\prime},\ldots,t_{m}^{\prime})+\ldots+g(0,t_{2}^{\prime},\ldots,t_{m}^{\prime}-1))L^{(0,t_{2}^{\prime},\ldots,t_{m}^{\prime})}=\sum_{t_{1}^{\prime}+\ldots+t_{m}^{\prime}=j\atop m=k-j}g(t_{1}^{\prime},t_{2}^{\prime},\ldots,t_{m}^{\prime})L^{(t_{1}^{\prime},\ldots,t_{m}^{\prime})}$